For their stats exam, the class average was 80.00 (sd= 11). Finn tells Jake his Z-score was 1.50
a. What was his score?
b. Convert his score to a distirbution where M= 100, SD= 10 ?
for a...I got
X= 80+11(1.50)
X= 80+ 16.50
X= 96.50
for b
I feel like im wrong in my calculation but i tried this:
96.50-100= -3.50/10= -0.35

Answer by stanbon(75887)

For their stats exam, the class average was 80.00 (sd= 11).
Finn tells Jake his Z-score was 1.50
a. What was his score?
b. Convert his score to a distirbution where M= 100, SD= 10 ?
for a...I got
X= 80+11(1.50)
X= 80+ 16.50
X= 96.50
Correct
-------------------
for b
If z-score = 1.5, X = 100 + 1.5*10 = 115
----------------
Cheers,
Stan H.